OpenStudy (anonymous):

Im getting confused with this:::: find the derivative of: r = 2/s^3 - 5/s

OpenStudy (anonymous):

\[the derivative = 2x ^{-3}-5x ^{-1}\]

OpenStudy (anonymous):

an then you do this to derive it....

OpenStudy (anonymous):

\[derivative= -6x+5x ^{-2}\]

OpenStudy (anonymous):

the above is your answer

OpenStudy (anonymous):

or you can write the answer like this...

OpenStudy (anonymous):

\[derivative = -6x+(5/x ^{2})\]

OpenStudy (anonymous):

Fan me if that helped :D

OpenStudy (anonymous):

why does the 5/s change to 5s^-1?? can you please tell me how to change

OpenStudy (anonymous):

There is a rule with derivatives that says that if there is for example 5/s then you can write it as well as 5s^-1

OpenStudy (anonymous):

other examples are 288/x^7 = 288x^-7

OpenStudy (anonymous):

IF 5/s if equal to 5s^-1 then why is it 5s^-2??

OpenStudy (anonymous):

actually, its the rules of exponents. So when you have \[X ^{-a} = 1/(X ^{a})\]

OpenStudy (anonymous):

the negative sign represents the inverse.

OpenStudy (anonymous):

What I did there was following that rule where you multiply the exponent by the coefficient. so what I did was get =2x^(−3)−5x^(−1) using the rule of exponents. Then I multiplied the 2 in front of the x by the x's exponent (2X(-3)) = -6 then I subtracted 1 from the exponent -> (-3-1)= -4

OpenStudy (anonymous):

wow I just realised that I made a mistake

OpenStudy (anonymous):

the final answer should be -6x^(-4)+5x^(-2)

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

yeah so like I was saying before you subtract 1 from the exponents

OpenStudy (anonymous):

the final answer in the ANSWERS SHEET IS \[-6/s^{4} + 5/s^{2}\]

OpenStudy (anonymous):

in replying to your question above, \[r=(2/s ^{3}) - (5/s)\] \[r= 2s ^{-3} - 5s ^{-1}\] Find the derivative \[r'= (-3)(2s ^{-3-1}) - (-1)(5s ^{-1-1})\] \[r' = -6s ^{-4} + 5s ^{-2}\]

OpenStudy (anonymous):

thank you heaps... that is great ~ i can actually see what is happening :)

OpenStudy (anonymous):

no problem. calculus is very easy if you just follow all the rules properly =)

OpenStudy (anonymous):

do u no a good website where i can learn calculus quickly?

OpenStudy (anonymous):

Calculus is actually something best learned in the setting of a classroom. But if you have any questions to ask just ask here and ill be more than happy to assist you.

OpenStudy (anonymous):

great thanks heaps :)

OpenStudy (anonymous):

yup